OpenStudy (anonymous):

Match each absolute value sentence with its solution set. Potential Matches: (1) {x|x > 4 or x < -4/5} (2) x < -1/3 or x > 1 (3) {4/3, -4} (4) {x| -11 < x < -7} (5) no solution Answer A : |x + 9| < 2 B: |5x - 8| > 12 C : |3k + 4| = 8 D : |x - 6| = -10 E : |-6x + 2| > 4

hartnn (hartnn):

do you know what absolute sign is ? and what is its effect ? like what is |-4| =... ?

OpenStudy (anonymous):

either -4 or 4

hartnn (hartnn):

no, its just 4 because |....| makes a negative number , positive. |-a| = a (if a>0) |-a| =-a (if a<0) got this ?

OpenStudy (anonymous):

uhh.. no. I got you up until you added the last line.

hartnn (hartnn):

|-a| =-a (if a<0) this ? let a=-1 then |-(-1)|= -(-1)=1 ok?

OpenStudy (anonymous):

okay..

hartnn (hartnn):

so, basically |....| can never be negative. if you find something like |.....| =-10 then there will be no solution to this.

OpenStudy (anonymous):

Okay. This is making a little more sense to me. Can you tell me how to actually figure one out though? Preferably one with a < or > sign?

hartnn (hartnn):

yes, i gave you that so that you can solve D. when \(|x|>a \implies x>a , \: x<-a \\or, -a<x<a\) similarly, \(|x|<a \implies x<a , \: x>-a \\or, -a>x>a\) so, what about \(|x+9|<2\) ?

OpenStudy (anonymous):

{x| -11 < x < -7}??

hartnn (hartnn):

just a second, i messed up with signs,sorry... yes, i gave you that so that you can solve D. when \(|x|>a \implies x>a , \: x<-a \\or, -a>x>a\) similarly, \(|x|<a \implies x<a , \: x>-a \\or, -a<x<a\) so, what about |x+9|<2 ? and yes, that is correct.!

OpenStudy (anonymous):

A4 B1 C3 D5 E2 Would these be right?

hartnn (hartnn):

for \(|x|=a \implies x=a, x=-a\) so, for |3k + 4| = 8, you get (3k+4)=8, (3k+4)=-8 can you solve these two normal equations to get 2 values of 'x' ?

hartnn (hartnn):

yes, those are correct! are you sure you got them after understanding the whole thing ?? ask if you have any sort of doubts....

OpenStudy (anonymous):

Yea I get it now. I have exams in a week and I'm studying and I couldn't remember how to do these to save my life! Thank you so much!

hartnn (hartnn):

ok, welcome ^_^